# Multimodal-enhanced Federated Recommendation: A Group-wise Fusion Approach

### v1

@@ Abstract.
Federated Recommendation (FR) is a new learning paradigm to tackle the learn-to-rank problem in a privacy-preservation manner. How to integrate multi-modality features into federated recommendation is still an open challenge in terms of efficiency, distribution heterogeneity, and fine-grained alignment. To address these challenges, we propose a novel multimodal fusion mechanism in federated recommendation settings (GFMFR). Specifically, it offloads multimodal representation learning to the server, which stores item content and employs a high-capacity encoder to generate expressive representations, alleviating client-side overhead. Moreover, a group-aware item representation fusion approach enables fine-grained knowledge sharing among similar users while retaining individual preferences. The proposed fusion loss could be simply plugged into any existing federated recommender systems empowering their capability by adding multi-modality features. Extensive experiments on five public benchmark datasets demonstrate that GFMFR consistently outperforms state-of-the-art multimodal FR baselines.
Federated Learning, Multimodal Recommender Systems
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: $X$; 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Driven by rising privacy concerns, Federated Recommendation (FR) has emerged as a prominent paradigm that enables collaborative model training without accessing raw user data, drawing increasing attention (Javeed et al., 2023; Sun et al., 2024; Wang et al., 2024). Despite this momentum, most FR methods rely heavily on structured item identifiers, such as unique IDs or tabular attributes, to model preferences. This narrow focus limits the ability to capture rich semantics from real-world items, which often include multimodal content like text, images, and audio(Liu et al., 2024a; Malitesta et al., 2025; Yang et al., 2025b). This substantially limits FR’s applicability, especially in domains where content signals play a central role in capturing user preferences(Liu et al., 2024b; Malitesta et al., 2024; Liu et al., 2023).
@@
The effective integration of multimodal information into FR remains underexplored, primarily due to several inherent challenges. First, multimodal item attributes often have large raw data sizes and require complex processing for semantic extraction. This imposes substantial storage and computation burdens on resource-constrained client devices. Second, heterogeneous data distributions and varying modality usage across clients make it difficult to learn representations that generalize well while preserving personalization. Third, aligning content features with recommendation objectives requires more than basic feature fusion, demanding explicit guidance to avoid redundancy and noise. These challenges highlight the need for a principled framework capable of systematically integrating multimodal knowledge into FR models.
@@
To this end, we introduce a conceptual framework named GFMFR (Group-wise Fusion for Multimodal-enhanced Federated Recommendation). The framework proposes a reallocation of learning responsibilities: Multimodal representation learning takes place centrally on the server side, where comprehensive item content and sufficient computational capacity are accessible. Clients then operate on compact, distilled representations that encapsulate both semantic content and user preference relevance. This design reduces client-side burdens and aligns well with commonly adopted system architectures in real-world deployments.
In addition to improving system efficiency, GFMFR incorporates a group-aware multimodal aggregation mechanism, which enables shared representation learning among users with similar preferences. This facilitates cross-client knowledge transfer while maintaining personalized modeling capabilities. Moreover, a preference-guided distillation strategy is introduced to ensure that multimodal signals directly contribute to the learning of recommendation-relevant features. Extensive experiments demonstrate that our method achieves state-of-the-art performance on benchmark datasets. Our main contributions are summarized as follows,
@@ 2.1. Multimodal Recommender System
With the proliferation of multimedia platforms and diverse content (e.g., short videos, images, news), incorporating multimodal information into recommender systems has become crucial for capturing user preferences and item characteristics. Such integration not only enriches representations but also alleviates data sparsity issues (Liu et al., 2024a; Zhang et al., 2024a; Zhao et al., 2025), making multimodal recommendation a prominent research direction (Ye et al., 2025; Zhuang et al., 2025; Zhou et al., 2025). The typical pipeline involves three stages: raw feature representation (Javed et al., 2021; Devlin et al., 2019), feature fusion (Wang et al., 2021; Tao et al., 2022; Yang et al., 2025a), and recommendation prediction (Koren et al., 2009; Meng et al., 2025). Raw feature extraction aims to obtain informative modality-specific representations (Chen et al., 2022; He et al., 2016); fusion methods (Zhou et al., 2023; Ma et al., 2022) integrate them into a unified space; and prediction models (He and McAuley, 2016; Xu et al., 2025) generate personalized recommendations. Despite their effectiveness, existing centralized approaches raise privacy concerns, limiting their applicability. To address this, we propose a multimodal recommendation framework under federated learning, ensuring privacy protection while retaining recommendation quality.
@@ 2.2. Federated Recommender System
Federated recommendation (FR) offers personalized services while safeguarding user privacy, making it well-suited for privacy-sensitive scenarios (Alamgir et al., 2022; Chronis et al., 2024; Harasic et al., 2024; Li and Long, 2024; Zhang et al., 2025b). Existing FR research spans two major perspectives: recommender systems and federated learning. From the recommender perspective, studies have explored diverse model architectures, including MF (Ammad-Ud-Din et al., 2019; Chai et al., 2020), MLP (Muhammad et al., 2020; Perifanis and Efraimidis, 2022), GNN (Wu et al., 2022; Zhang et al., 2024b), and Transformer (Wei et al., 2023; Zhang et al., 2025a), as well as application scenarios such as cross-domain (Meihan et al., 2022; Zhang et al., 2024d), cold-start (Zhang et al., 2024c; Li et al., 2025), and news recommendation (Qi et al., 2020; Yi et al., 2021). From the federated learning perspective, efforts have addressed core challenges in security (Liu et al., 2022; Perifanis et al., 2023), robustness (Zhang et al., 2022; Ali et al., 2024), and efficiency (Zhang et al., 2023b; Nguyen et al., 2024). However, most FR methods still rely on ID-based item representations, overlooking rich multimodal content (e.g., text, images), which can offer valuable semantic cues.
@@
Research on multimodal FR remains limited. Existing methods such as AMMFRS (Feng et al., 2024) and FedMR (Li et al., 2024c) adopt client-side fusion strategies, integrating multimodal features locally: AMMFRS via attention-based weighting and FedMR via interaction-aware fusion modules. However, these approaches assume clients can store and process large-scale multimodal data, which is impractical for resource-constrained devices. FedMMR (Li et al., 2024a) shifts fusion to the server and aligns modalities using a contrastive loss, but focuses primarily on cross-modal alignment while overlooking inter-modal semantic dependencies. To address these limitations, we propose a novel server-side multimodal fusion framework that offloads both storage and representation learning. By further incorporating inter-user correlations, our approach integrates multimodal signals in a fine-grained manner, improving recommendation quality while maintaining efficiency.
3. Problem Formulation
In this section, we present a formalization of our proposed conceptual framework for integrating multimodal content into federated recommender systems. We begin with the general formulation of FR, followed by a clear clarification of the problem and a detailed explanation of the corresponding optimization objective.
@@ 3. Problem Formulation
General FR framework. Let $\mathcal{U}$ and $\mathcal{I}$ denote the user set and item set in the recommender system, and $N=|\mathcal{U}|$ and $M=|\mathcal{I}|$ are the total number of users and items respectively. For a recommendation model $\mathcal{F}$ parameterized by $\theta$, it is generally composed of three modules, i.e., user embedding $E_{\mathcal{U}}$, item embedding $E_{\mathcal{I}}$ and prediction function $P$. Given user $u$ and item $i$, the recommendation model makes rating predictions as $\hat{\mathcal{R}}_{ui}=\mathcal{F}(u,i|\theta)$. In the federated recommender system, each user $u$ acts as a client who holds the personal data $\mathcal{R}_{u}$, including all the historically rated items. To implement an FR model, the system alternates between the clients and server and performs the following steps:
@@
Figure 1. Overview of the proposed GFMFR. The left part illustrates a general FR framework. Given $k$ modalities of raw input data $\{X_{m}\}_{m=1}^{k}$, GFMFR builds upon this by introducing an additional fusion loss $\mathcal{L}_{fus}$, which integrates the ID-based feature $E_{\mathcal{I}}$ learned from the recommendation model $\mathcal{F}_{\theta}$ with multimodal feature $E_{\mathcal{M}}$ learned by a multimodal representation model $\mathcal{H}_{\psi}$, thereby enabling multimodal information integration into the federated recommendation system.
The proposed multimodal-enhanced FR framework. Given raw multimodal item data $\{X_{m}\}_{m=1}^{k}$, where $k$ is the number of item modalities. Let $\mathcal{H}_{\psi}$ denotes the multimodal representation learning model, and $E_{\mathcal{M}}=\mathcal{H}_{\psi}(\{X_{m}\}_{m=1}^{k})$ denotes the learned multimodal representation. The multimodal-enhanced FR framework aims to integrate the multimodal representation $E_{\mathcal{M}}$ into the item embedding $E_{\mathcal{I}}$ through a dedicated fusion loss $L_{fus}$, thereby improving the overall effectiveness of FR learning. Formally, we formulate the optimization objective of the multimodal-enhanced FR framework as,
(2) $\min_{\{E_{\mathcal{U}}^{u}\}_{u=1}^{N},E_{\mathcal{I}},P}\sum_{u=1}^{N}\alpha_{u}(\mathcal{L}_{rec}(\theta;\mathcal{R}_{u})+\lambda\cdot\mathcal{L}_{fus}(E_{\mathcal{I}},E_{\mathcal{M}}))$
where $\lambda$ is the coefficient of fusion loss. The proposed multimodal-enhanced FR framework is inherently general and compatible, allowing for straightforward integration into existing FR models with minimal architectural modifications. The framework overview is illustrated in Figure 1.
@@ 4.1. Multimodal Representation Initialization
In multimodal recommendation, item profiles comprise diverse modalities such as text, images, and audio. To leverage these signals, raw inputs from each modality are encoded into model-compatible representations via modality-specific encoders. The server employs a powerful multimodal representation model to produce initial embeddings $\mathcal{E}_{m}\in\mathbb{R}^{M\times d_{1}}$ for $m=1,\ldots,k$, where $d_{1}$ denotes the embedding dimension. Leveraging abundant server resources, large-scale pre-trained models serve as backbones to generate high-quality initial representations.
@@ 4.2. Group-Aware Multimodal Aggregation
In FR systems, heterogeneous data distributions and modality usage across clients cause inconsistent multimodal representations, impeding knowledge sharing. Global representation enforcement improves consistency but compromises personalization. To address this, we propose a group-aware multimodal aggregation that derives shared representations for preference-aligned user groups, enabling efficient knowledge transfer while preserving client-specific modeling.
Cluster clients into groups. Clients are clustered by their prediction functions, which capture decision patterns over items and group users with similar preferences. This method preserves privacy by eliminating the need to access raw user data. Specifically, the server aggregates clients’ prediction functions $\{P_{u}\}_{u=1}^{N}$ and clusters them based on their parameters, formally expressed as,
@@
where $g$ denotes the total number of user groups, and common choices for the clustering algorithm include K-Means and hierarchical clustering.
Learn group-level multimodal features. Given initial representations $\{\mathcal{E}_{m}\}_{m=1}^{k}$ and user groups $\{G_{1},...,G_{g}\}$, a multimodal aggregation module learns shared features within each group. This module comprises a learnable group embedding $E_{G}\in\mathbb{R}^{g\times d_{2}}$ assigning unique IDs to groups, and an aggregation function $S_{\phi}$. For each group $l=1,\dots,g$, the aggregation is formulated as,
@@
where $E_{G}(l)$ is the embedding for group $l$, and $Q_{l}$ denotes the aggregated multimodal features. To optimize this module, we use group-level item embeddings as supervision. Specifically, for group $l$, the item embedding $C_{l}$ is obtained by aggregating local embeddings from members, e.g., $C_{l}=\text{AvgPooling}(\{E_{\mathcal{I}}^{u}\}_{u\in G_{l}})$. The parameters of $S_{\phi}$ are then updated by minimizing,
@@
where $\mathcal{L}_{agg}$ is the aggregation loss, e.g., the mean squared error between $Q_{l}$ and $C_{l}$. This optimization aligns aggregated group-level multimodal features with group-specific item embeddings, enhancing cross-modal consistency and yielding representations better suited for downstream recommendation.

### v2

# arXiv:2509.19955v2[cs.IR] 27 Feb 2026

@@ Abstract.
Federated Recommendation (FR) has emerged as a promising paradigm for addressing the learn-to-rank problem in a privacy-preserving manner. However, effectively incorporating multimodal item features into FR remains an open challenge, due to efficiency constraints, distribution heterogeneity, and feature utilization alignment with the recommendation objective. To tackle these issues, we propose GFMFR, a novel multimodal fusion framework for federated recommendation. Specifically, multimodal representation learning is offloaded to the server, which stores item content and employs a high-capacity encoder to generate expressive representations, thereby alleviating the computational burden on clients. In addition, a group-aware multimodal aggregation mechanism learns shared representations for users with similar interests, enabling knowledge sharing while alleviating distribution heterogeneity. Finally, GFMFR adopts a preference-guided distillation strategy that leverages multimodal information in a way directly aligned with recommendation objectives. The proposed framework can be seamlessly integrated into existing federated recommender systems, enhancing their effectiveness by incorporating multimodal features. Extensive experiments on five benchmark datasets demonstrate that GFMFR consistently outperforms state-of-the-art multimodal FR baselines. The implementation code is available [^1] https://github.com/Zhangwp2420/GFMFR.
Federated Learning, Multimodal Recommender Systems, User Personalization Modeling
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792573ccs: Information systems Personalizationccs: Information systems Recommender systems
@@ 1. Introduction
Driven by rising privacy concerns, Federated Recommendation (FR) has emerged as a prominent paradigm that enables collaborative model training without accessing raw user data, drawing increasing attention (Javeed et al., 2023; Sun et al., 2024; Wang et al., 2024). Despite this momentum, most existing FR methods still depend heavily on structured item identifiers, such as unique IDs or tabular attributes, to represent user preferences. This limitation becomes particularly pronounced in web-based commerce platforms, where user–item interactions are mediated by multimodal product content such as textual descriptions and images (Liu et al., 2024a; Malitesta et al., 2025; Yang et al., 2025b). Product multimodality is indispensable for capturing fine-grained semantics of items and reflecting user intent at scale, yet it has received limited attention in the FR literature. This gap substantially weakens the applicability of current FR models to realistic Web scenarios (Liu et al., 2024b; Malitesta et al., 2024; Liu et al., 2023).
@@
However, the effective integration of multimodal information into FR remains underexplored, primarily due to several inherent challenges. First, multimodal item attributes often have large raw data sizes and require complex processing for semantic extraction. This imposes substantial storage and computation burdens on resource-constrained client devices. Second, heterogeneous data distributions, particularly differences in modality preferences across clients, make it difficult to learn a unified multimodal representations that generalize well while preserving personalization. Third, aligning multimodal features with recommendation objectives remains a difficult task. Basic fusion methods often introduce redundancy and noise, making it challenging to preserve semantics that are directly relevant for modeling user preferences. These considerations highlight the need for a systematic framework to exploit multimodal knowledge in FR models.
@@
To this end, we introduce a new framework, Group-wise Fusion for Multimodal-enhanced Federated Recommendation (GFMFR). The framework proposes a reallocation of learning responsibilities: Multimodal representation learning takes place centrally on the server side, where comprehensive item content and sufficient computational capacity are accessible. Clients then operate on compact, distilled representations that encapsulate both semantic content and user preference relevance. This design reduces client-side burdens and aligns well with commonly adopted system architectures in real-world deployments (Solution for challenge 1). In addition to improving system efficiency, GFMFR incorporates a group-aware multimodal aggregation mechanism, which enables shared multimodal representation among users in a group with similar preferences. By coordinating learning across users within a group, the mechanism mitigates cross-client heterogeneity while enabling collaborative knowledge propagation (Solution for challenge 2). Moreover, a preference-guided distillation strategy is introduced. It maps multimodal representations into the user preference space and incorporates them into client model optimization via a distillation loss. This approach ensures that multimodal information is utilized in a way that directly aligns with the recommendation objective (Solution for challenge 3). Extensive experiments demonstrate that our method achieves state-of-the-art performance on multimodal recommendation benchmark datasets. Our main contributions are summarized as follows,
@@ 2.1. Multimodal Recommender System
With the proliferation of multimedia platforms and diverse content (e.g., short videos, images, news), incorporating multimodal information into recommender systems has become crucial for capturing user preferences and item characteristics. Such integration not only enriches representations but also alleviates data sparsity issues (Liu et al., 2024a; Zhang et al., 2024d; Zhao et al., 2025), making multimodal recommendation a prominent research direction (Ye et al., 2025; Zhuang et al., 2025; Zhou et al., 2025). The typical pipeline involves three stages: raw feature representation (Javed et al., 2021; Devlin et al., 2019), feature fusion (Wang et al., 2021; Tao et al., 2022; Yang et al., 2025a), and recommendation prediction (Koren et al., 2009; Meng et al., 2025). Raw feature extraction aims to obtain informative modality-specific representations (Chen et al., 2022; He et al., 2016); fusion methods (Zhou et al., 2023; Ma et al., 2022) integrate them into a unified space; and prediction models (He and McAuley, 2016; Xu et al., 2025) generate personalized recommendations. Despite their effectiveness, existing centralized approaches raise privacy concerns, limiting their applicability. To address this, we propose a multimodal recommendation framework under federated learning, ensuring privacy protection while retaining recommendation quality.
@@ 2.2. Federated Recommender System
Federated recommendation (FR) offers personalized services while safeguarding user privacy, making it well-suited for privacy-sensitive scenarios (Alamgir et al., 2022; Chronis et al., 2024; Harasic et al., 2024; Li and Long, 2024; Zhang et al., 2025b). Existing FR research spans two major perspectives: recommender systems and federated learning. From the recommender perspective, studies have explored diverse model architectures, including MF (Ammad-Ud-Din et al., 2019; Chai et al., 2020), MLP (Muhammad et al., 2020; Perifanis and Efraimidis, 2022), GNN (Wu et al., 2022; Zhang et al., 2024a), and Transformer (Wei et al., 2023; Zhang et al., 2025a), as well as application scenarios such as cross-domain (Meihan et al., 2022; Zhang et al., 2024c), cold-start (Zhang et al., 2024b; Li et al., 2025), and news recommendation (Qi et al., 2020; Yi et al., 2021). From the federated learning perspective, efforts have addressed core challenges in security (Liu et al., 2022; Perifanis et al., 2023), robustness (Zhang et al., 2022; Ali et al., 2024), and efficiency (Zhang et al., 2023b; Nguyen et al., 2024). However, most FR methods still rely on ID-based item representations, overlooking rich multimodal content (e.g., text, images), which can offer valuable semantic cues.
@@
Research on multimodal FR remains limited. Existing methods such as AMMFRS (Feng et al., 2024) and FedMR (Li et al., 2024c) adopt client-side fusion strategies, integrating multimodal features locally: AMMFRS via attention-based weighting and FedMR via interaction-aware fusion modules. However, these approaches assume clients can store and process large-scale multimodal data, which is impractical for resource-constrained devices. FedMMR (Li et al., 2024a) shifts fusion to the server and aligns modalities using a contrastive loss, but focuses primarily on cross-modal alignment while overlooking inter-modal semantic dependencies. To address these limitations, we propose a novel server-side multimodal fusion framework that offloads both storage and representation learning. By performing semantic fusion across modalities and further leveraging collaboration among users with similar preferences, our approach enables more coherent representation learning and fosters collaborative knowledge sharing in FR systems.
3. Preliminary
General FR framework.
@@ 3. Preliminary
Let $\mathcal{U}$ and $\mathcal{I}$ denote the user set and item set in the recommender system, and $N=|\mathcal{U}|$ and $M=|\mathcal{I}|$ are the total number of users and items respectively. For a recommendation model $\mathcal{F}$ parameterized by $\theta$, it is generally composed of three modules, i.e., user embedding $E_{\mathcal{U}}$, item embedding $E_{\mathcal{I}}$ and prediction function $P$. Given user $u$ and item $i$, the recommendation model makes rating predictions as $\hat{\mathcal{R}}_{ui}=\mathcal{F}(u,i|\theta)$. In the federated recommender system, each user $u$ acts as a client who holds the personal data $\mathcal{R}_{u}$, including all the historically rated items. To implement a FR model, the system alternates between the clients and server and performs the following steps:
@@ 4. Group-wise Fusion for Multimodal-enhanced FR (GFMFR)
Figure 1. Overview of the proposed GFMFR. The entire process unfolds in five stages. Initially, before the federated optimization begins, the server generates initial multimodal representations based on raw multimodal data (Step ①). During each communication round, clients first train local recommendation models using their private data and upload the resulting item embeddings and prediction functions to the server (Step ②). The server then performs global aggregation on the received embeddings and functions, groups users based on their prediction functions, and computes group-level item embeddings and prediction functions (Step ③). Leveraging these group-level embeddings, the server trains a multimodal aggregation model and projects the fused representations into the preference space via the group-level prediction functions (Step ④). Finally, the global parameters and group-level preference representations are sent back to the clients to refine their local models (Step ⑤).
@@ 4.1. Multimodal Representation Initialization
In multimodal recommendation, item profiles are enriched with various signals such as textual descriptions and product images. To make these diverse sources usable, raw inputs from each modality should first be processed by modality-specific encoders, ensuring that their distinct characteristics are preserved in model-compatible form. Handling these inputs on clients is often infeasible due to limited computational and storage resources, whereas the server is equipped with abundant computational capacity and can leverage large-scale pre-trained backbones to generate semantically rich and well-structured representations. To utilize these signals effectively, this work relies on the server to learn initial embeddings directly from the raw data using a multimodal representation learning model, producing initial embeddings $\mathcal{E}_{m}\in\mathbb{R}^{M\times d_{1}}$ for $m=1,\ldots,k$, where $k$ is the number of item modalities and $d_{1}$ denotes the embedding dimension, thereby forming a solid foundation for subsequent multimodal fusion and integration of the representations in FR systems.
@@ 4.2. Group-Aware Multimodal Aggregation
In FR systems, users exhibit substantial heterogeneity in how they rely on different modalities. For instance, younger users may focus more on visual cues, whereas older users may place greater emphasis on textual descriptions. This diversity means that a single global multimodal representation cannot accommodate all users, while learning fully user-specific fusion quickly becomes inefficient and unstable. To better balance generalization and personalization, we propose a group-aware multimodal aggregation mechanism that learns shared representations for preference-aligned user groups, capturing intra-group regularities and enabling more effective multimodal knowledge transfer.
4.2.1. Cluster clients into groups.
To effectively capture shared patterns among users with similar preferences, we begin by organizing clients into groups, which forms the basis for learning group-level representations. This grouping is based on clients’ prediction functions, which reflect how users make decisions over items and thus implicitly encode their preference patterns. Instead of directly sharing raw interaction or feature data, each client transmits its learned prediction function $\{P_{u}\}_{u=1}^{N}$, preserving privacy while still providing meaningful signals for grouping. The server then collects the set of functions $\{P_{u}\}_{u=1}^{N}$ and applies a clustering algorithm to partition users into groups with similar decision behaviors, formally expressed as
@@
where $g$ denotes the total number of user groups. Standard clustering methods such as K-Means or hierarchical clustering can be applied, and the resulting groups serve as the basis for learning shared group-level multimodal representations.
4.2.2. Learn group-level multimodal features.
Once the user groups $\{G_{1},...,G_{g}\}$ have been identified, we learn shared multimodal representations within each group. Given the initial item representations $\{\mathcal{E}_{m}\}_{m=1}^{k}$ produced by the server, a multimodal aggregation module integrates these features for each group, producing group-level embeddings that consolidate the multimodal signals of items with respect to the user cluster.
The aggregation module consists of a learnable group embedding $E_{G}\in\mathbb{R}^{g\times d_{2}}$, which assigns a unique vector representation to each user group, and an aggregation function $S_{\phi}$ that combines item-level multimodal embeddings with the group embedding. Formally, for each group $l=1,\dots,g$, the aggregation is formulated as,
@@
where $E_{G}(l)$ is the embedding for group $l$, and $Q_{l}$ denotes the aggregated multimodal features.
To optimize this module, we leverage group-level item embeddings as supervisory signals, guiding the aggregated multimodal features to better reflect the collective item representations within each user group. Specifically, for each group $l$, the group item embedding $C_{l}$ is obtained by aggregating the local embeddings of its members, e.g., $C_{l}=\text{AvgPooling}(\{E_{\mathcal{I}}^{u}\}_{u\in G_{l}})$. These embeddings summarize how items are perceived by users within the group, providing a target for group-level aggregation. The parameters of the aggregation function $S_{\phi}$ and the group embeddings $E_{G}$ are then updated by minimizing,
@@
where $\mathcal{L}_{agg}$ denotes the aggregation loss, typically instantiated as the mean squared error between the aggregated group embedding $Q_{l}$ and the corresponding group item embedding $C_{l}$. This optimization encourages the group-level multimodal representations to capture consistent cross-modal semantics while reflecting the collective perception of items within each user group, facilitating more coherent and informative embeddings for subsequent recommendation learning.